OpenStudy (666potatoes):

Rationalize the denominator. 4/(3 + 5i) I already multiplied denominator and numerator by 3+5i and got 12+20i/16+30i. Now what? :l

OpenStudy (anonymous):

\[{4 \over (3 + 5i)}\times {(3-5i) \over (3-5i)}\]You need to multiply it with a switched sign. Try this

OpenStudy (anonymous):

\[{12-20i \over 34}=\frac{ 6-10i }{ 17 }\]

OpenStudy (666potatoes):

Ty ty. c: